Three point charges are arranged on a line. Charge q3= +5.00 nC and is at the origin. Charge q2= -3.00 nC and is at x= +4.00 cm. Charge q1 is at x= +2.00 cm. What is q1 (magnitude and sign) if the net force on q3 is zero?


1
Expert's answer
2021-04-05T11:10:16-0400

The distance between "q_3" and "q_2" is "r_{32} = 4m", the distance between "q_3" and "q_1" is "r_{31} = 2m".

According to the Coulomb's law, the force on "q_3" from "q_2" is:


"F_{32} =k \dfrac{q_3q_2}{r_{32}^2}"

where "k\approx 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant.

The force on "q_3" from "q_1" is:


"F_{31} =k \dfrac{q_3q_1}{r_{31}^2}"

Since these forces are directed along one line, we can add them. The resultant force is then:


"F = F_{32} + F_{31} = 0\\
k \dfrac{q_3q_1}{r_{31}^2} = -k \dfrac{q_3q_2}{r_{32}^2} \\
\dfrac{q_1}{r_{31}^2} = -\dfrac{q_2}{r_{32}^2}\\"

Finally:


"q_1 = -q_2\dfrac{r_{31}^2}{r_{32}^2}"


Substituting the numbers, obtain:


"q_1 = -(-3\times 10^{-9})\dfrac{2^2}{4^2}=+ 0.75\times 10^{-9}C = +0.75\space nC"

Answer. "+0.75\space nC".